OpenStudy (anonymous):

Express in scientific notation. (6*10^a)(5*10^b)

OpenStudy (lgbasallote):

hint: \[\huge (m \times 10^x)(n \times 10^y) \implies mn \times 10^{x+y}\] does that help?

OpenStudy (anonymous):

30*10^ab

OpenStudy (lgbasallote):

\[30 \times 10^{a+b}\] you add the exponents

OpenStudy (anonymous):

oh. srry